Where can one get a deck of personalized playing cards?
A deck of personalized playing cards can be found on sites like Shutterfly, Printer Studio, Walgreens Photo, Shindigz, Artscow, Finn Custom, and Ebay.
Can the probability be more than 1 or less than 0?
No, all probabilities are given from a scale of 0 to 1. 0 represents 0% and 1 represents a 100% probability.
What is the probability to get face cards if a card is drawn from a pack of 52?
Counting Aces as face cards, the probability is 16/52 or 4/13.
vulnerability
What are equally likely event?
Two events that have the same chance of happening. For example, if I flip a coin the event of obtaining a 'head' is equally as likely as the event of obtaining a 'tail'.
But equally likely does not mean 0.5 probability. It's possible that it's equally likely that someone in Ontario, Canada will die from being stung by a wasp as from being electrocuted in their kitchen at home. Neither event is very likely but the two events could be equally likely.
What is the probability of rolling a multiple of 3?
The answer depends on how often the die is rolled. As the number of rolls increases, the probability gets near enough to 1 as makes no difference.
For a single roll, the answer is 1/3.
What is an example of relative frequency probability?
Flip a coin 1000 times, counting the number of 'heads' that occur. The relative frequency probability of 'heads' for that coin (aka the empirical probability) would be the count of heads divided by 1000. Please see the link.
If a single die is rolled 9 times what is the probability of getting exactly 2 sixes?
There are 69 total possibilities.
Imagine having nine boxes, that each hold how many possibilities each individual dice can hold. Any two of the dice can only be a six, therefore any two of the boxes can only have a 1 (only one possibility). The rest of the boxes can be anything, other than 6 (so the rest of the boxes have 5 possibilities each).
[1] [1] [5] [5] [5] [5] [5] [5] [5]
Now multiply all the possibilities together you get total number of combinations. In this case, it's 57. You can use this same strategy to see why the total number of possibilities is 69.
The answer is total restricted combinations divided by total combinations.
57 / 69 ~= 0.007752 (or roughly 0.78%)
120 is one standard deviation greater than the mean (z = 1).
So you want Pr(z < 1)
The probability of a score at most 1 sd away from the mean is 0.68
That is Pr(-1 < z < 1) = 0.68
So Pr(|z| > 1) = 1 - 0.68 = 0.32
then by symmetry,
Pr(z > 1) = 1/2*0.32 = 0.16
So Pr(z < 1) = 1 - 0.16 = 0.84
120 is one standard deviation greater than the mean (z = 1).
So you want Pr(z < 1)
The probability of a score at most 1 sd away from the mean is 0.68
That is Pr(-1 < z < 1) = 0.68
So Pr(|z| > 1) = 1 - 0.68 = 0.32
then by symmetry,
Pr(z > 1) = 1/2*0.32 = 0.16
So Pr(z < 1) = 1 - 0.16 = 0.84
120 is one standard deviation greater than the mean (z = 1).
So you want Pr(z < 1)
The probability of a score at most 1 sd away from the mean is 0.68
That is Pr(-1 < z < 1) = 0.68
So Pr(|z| > 1) = 1 - 0.68 = 0.32
then by symmetry,
Pr(z > 1) = 1/2*0.32 = 0.16
So Pr(z < 1) = 1 - 0.16 = 0.84
120 is one standard deviation greater than the mean (z = 1).
So you want Pr(z < 1)
The probability of a score at most 1 sd away from the mean is 0.68
That is Pr(-1 < z < 1) = 0.68
So Pr(|z| > 1) = 1 - 0.68 = 0.32
then by symmetry,
Pr(z > 1) = 1/2*0.32 = 0.16
So Pr(z < 1) = 1 - 0.16 = 0.84
There are not just 4 probabilities. Probability is a continuous variable ranging from 0 to 1: it can take infinitely many values.
There are not just 4 probabilities. Probability is a continuous variable ranging from 0 to 1: it can take infinitely many values.
There are not just 4 probabilities. Probability is a continuous variable ranging from 0 to 1: it can take infinitely many values.
There are not just 4 probabilities. Probability is a continuous variable ranging from 0 to 1: it can take infinitely many values.
What can you do to get the experimental probability to be closer to the theoretical probability?
The experimental probability is figured out when a person goes through the trouble of actually trying it out. Theoretical probability is when a person comes to a conclusion of what is most likely, based off of the experiment results.
384,475,000 license plates.
There are 35 different letter/number combinations possible. Each combination has 10,985,000 variants. 35*10,985,000 = 384,475,000
What does it mean if the standard deviation is gr?
The answer will depend on what "gr" is.
The answer will depend on what "gr" is.
The answer will depend on what "gr" is.
The answer will depend on what "gr" is.
What is the probability of rolling a dice and getting an odd number?
The probability of getting an odd number in a single throw of a fair die (not dice!) is 1/2.
The probability of getting an odd number in a single throw of a fair die (not dice!) is 1/2.
The probability of getting an odd number in a single throw of a fair die (not dice!) is 1/2.
The probability of getting an odd number in a single throw of a fair die (not dice!) is 1/2.
How many square number cards in a deck?
8 four 4's and four 9's
unless you want to count aces as 1 then there would be 12
What are chances of 4 aces being in a hand of 9 cards?
The probability of 4 aces being in a hand of 9 cards is:
9C4 ∙ (4/52)∙(3/51)∙(2/50)∙(1/49)∙(48/48)∙(47/47)∙∙∙(44/44) = 0.0004654...
≈ 0.0465%
where 9C4 = 9!/[(9-3)!∙3!] = 126
What is the probability of rolling an odd number less than 11?
The answer depends on how many dice are rolled. With 12 or more dice, the answer is 0.
The answer depends on how many dice are rolled. With 12 or more dice, the answer is 0.
The answer depends on how many dice are rolled. With 12 or more dice, the answer is 0.
The answer depends on how many dice are rolled. With 12 or more dice, the answer is 0.
What is the probability found by experimenting?
It is experimental or empirical probability.
It is experimental or empirical probability.
It is experimental or empirical probability.
It is experimental or empirical probability.
We would need to know which events you're referring to to respond to this question.
What is the probability of not spinning a 4 with one spin?
The answer depends on the shape of the spinner.
This is the same as the probability of choosing either a red of a blue marble.
There are 5+4 out of 15 ways of doing this. The probability is therefore 9/15 = 3/5.